OpenStudy (cj49):

\[If \tan \theta = \sqrt{1-\cos2 \theta \div1+\cos2\theta}, then \tan (22 1/2 )degress=\]

OpenStudy (cj49):

\[If \tan \theta = \sqrt{1-\cos2 \theta \div1+\cos2\theta}, then \tan (22 1/2 )degress=\]

OpenStudy (cj49):

its 22 1/2

OpenStudy (solomonzelman):

you can use ~ as a space. If you put ~~ or ~~~ you will get more and more space. (meaning wider space)

OpenStudy (cj49):

thnx..